OpenStudy (anonymous):

Water is leaking out of a container. At 2:30 p.m., there are 39 pints of water in the container. At 4 p.m., there are 27 pints of water in the container. If the water continues leaking out of the container at the same rate, how much water will be in the container at 6:30 p.m.? 23 pints 47 pints 22 pints 7 pints

OpenStudy (anonymous):

I'm sorry, I'm bad at math :(

OpenStudy (anonymous):

its alright

OpenStudy (anonymous):

D.7 pints

OpenStudy (anonymous):

39-27=12 which mean it leak 12pints since 2:30 to 4, 90minutes and the water leak in one minute is 12/90 pints=2/15 from 4pm to 6:30 pm have 150minutes so the water leak between that time is 150*2/15=20pints 27-20=7 so answer is D